origination Fees.  Jack needs to borrow ​$500 for one year. Bank South will give him the loan at 11​%. SunCoast Bank will give him the loan at 8​% with a $23. loan origination fee. First National will give him the loan at 6% with a ​$9 loan origination fee. Determine the total interest and fees Jack will be charged in each case. Which loan should Jack​ choose?

* so i know the answer is $55 bank south.

than they asked *If Jack takes the loan at SunCoast​ Bank, the amount of total interest and fees is ? so i know is $63.

but I dont know how to get the $63. can anyone explained me. ???

please.
Answer #1

Bank south will charge interest of 11℅ so total interest = 500*11℅ = 55

Suncoast bank will charge interest of 8℅ so total interest = 500*8℅ = 40 plus it will charge loan origination fees of 23

So interest + fees = 40 + 23 = 63

First national will charge interest of 6℅ so total interest = 500*6℅ = 30 plus loan origination fees of 9

So interest + fees = 30 + 9 = 39

Hence jack should opt for first national